Attachment 1: The PD was removed from the transmission side of the OMC #002 (former LHO OMC - the one blasted by the optical pulse in Aug 2016).
It was confirmed that the PD has the scribing mark saying "A".
Attachment 2: This diode had no glass cap on it. The photodiode sensitive element is still intact. For ease of handling, it should be kept in a cage. There are four cages in the OMC lab, but they are ocuppied with the High QE PDs and others. So, the cage for this PD was offered by Rich from his office, meaning the cage was not clean.
Attachment 3: The sensor side is capped by a plate. This cap can be removed by unscrewing the two cap screws in the photo.
Attachment 4: The PD legs are shorted. (Just to match the style with the LLO one).
Attachment 5: Wrapped with AL foil and double bagged. (Repeat: It is not anything clean.)
Attachment 6: The bag was left on Rich's desk.